# Router-Tuning: A Simple and Effective Approach for Dynamic Depth

### v2

@@ 6 Ablation Studies
Table 4: Performance of Router-Tuning on Mixture of Experts, where we take Expert Drop He et al. (2024a) as the baseline of static dropping for comparison.
@@
Since Router-Tuning can be seamlessly integrated into pretrained language models, we extend our evaluation to other mainstream models, specifically Llama-2 Touvron et al. (2023), Mistral Jiang et al. (2023) and Qwen2.5 Bai et al. (2023), covering a wide range of models with varying sizes. In Table 3, we experiment with half of attention layers deployed with Router-Tuning and keep the total MoD capacity at 50%. Compared to the original models, the involvement of Router-Tuning maintains the performance in all the models, which demonstrates the effectiveness of Router-Tuning across different models.
Figure 4: Ablation study on the number of MoD layers.
@@
Impact of number of MoD layers
In the main experiments, we deployed half of the layers with MoD. In Figure 4, we further explore the effect of the number of MoD layers. Our results indicate that applying MoD to half of the attention layers maintains comparable performance. However, when further increasing the number of MoD layers, performance starts to degrade. We attribute this decline to the transformation of important shallow layers, which negatively impacts overall performance Men et al. (2024); He et al. (2024a, b). Therefore, preserving the density of shallow layers while applying MoD to deeper layers ensures its effectiveness.
Effect of training dataset
Since Router-Tuning requires a dataset to fine-tune the routers, we next examine the impact of using different training datasets in Table 5. We consider a variety of datasets, including Alpaca Taori et al. (2023), Evol-Instruct Xu et al. (2023), ShareGPT Zheng et al. (2023), and Llama-Pro Wu et al. (2024). Since Router-Tuning only fine-tunes the routers while keeping the backbone of the language models intact, changes in the training dataset do not significantly impact performance. However, Llama-Pro, which incorporates diverse training data from various domains, provides slightly better performance due to its broader data coverage. On the other hand, due to the small number of trainable parameters, Router-Tuning does not require a large amount of training samples. As shown in Figure 5, MoD layers are initially dense-activated and then sparsified. Although the initial sparsification steps lead to a drop in performance, subsequent Router-Tuning facilitates performance recovery. Notably, just 5K training samples are sufficient to effectively train the routers.
Figure 5: Ablation study on the number of training samples.
7 MoE and LoRA Integration
In this section, we further explore the integration of Router-Tuning with other architectures and training techniques. First, we implement Router-Tuning on mainstream MoE architectures. Then, we combine Router-Tuning with LoRA fine-tuning to enhance both efficiency and performance.
Table 6: Effectiveness of Router-Tuning integrated with LoRA finetuning, compared to deploying Router-Tuning and LoRA seperately.
@@ 7 MoE and LoRA Integration
The Expert’s redundancy in MoE has been widely demonstrated in recent works Lu et al. (2024); He et al. (2024a), e.g., the model still maintains comparable performance after removing certain layers. Therefore, we further extend Router-Tuning to MoE, where we take OLMoE Muennighoff et al. (2024) and DeepSeek-MoE Dai et al. (2024) as the backbones and equip each Expert network with Router-Tuning. Since these models deploy MoE at the token level, where each Expert processes tokens from different sequences within a batch, we directly apply token-level Router-Tuning to these models. Here, we compare Router-Tuning with Expert Drop He et al. (2024a) that drops unimportant experts. During deployment, we configure an overall skipping ratio of 25%, achieved either by statically dropping 25% of layers or utilizing MoD on half of the MoE layers with 50% MoD capacity.
Figure 6: Expert-wise MoD capacity, visualizing the capacity utilization of individual experts within a specific layer.
In Table 4, while Expert Drop affects the performance significantly, Router-Tuning preserves more than 95% of the original performance. In Figure 6, we visualize the capacity utilization of individual experts within a specific layer, considering two perspectives for each expert: the frequency of expert selection (“Frequency”) and the MoD capacity allocated to each expert (“Capacity”).
Rather than removing a subset of experts—typically those selected less frequently Router-Tuning adopts a specialized and efficient utilization pattern, where high-frequency experts undergo greater sparsification compared to low-frequency experts. This approach retains the potential of all experts and fosters a more balanced utilization across the entire set of experts Fedus et al. (2022). Therefore, this approach maintains the potential of all experts and promotes a more balanced utilization across all experts.
@@ 8 Conclusion
Despite the progress we have made, our work still has limitations. First, while we have advanced MoD with Router-Tuning, other sophisticated training methods may exist that could further improve performance, warranting future exploration. Second, due to computational resource constraints, our experiments have been limited to Llama-3-8B, Mistral-7B, and Qwen2.5-7B on a small set of tasks. Expanding this approach to other models and a broader range of tasks would be highly valuable for understanding its full potential.

### v4

@@ 6 Ablation Studies
We conduct ablation studies to evaluate the generality and robustness of Router-Tuning across various settings, including different model architectures, the number of MoD layers, training datasets.
Table 4: Performance of Router-Tuning on Mixture of Experts, where we take Expert Drop He et al. (2024a) that statically skips the experts as the baseline of static dropping for comparison.
@@
Since Router-Tuning can be seamlessly integrated into pretrained language models, we extend our evaluation to diverse architectures, including Llama-2 Touvron et al. (2023), Mistral Jiang et al. (2023), and Qwen2.5 Bai et al. (2023), covering a wide range of model sizes. As shown in Table 3, we deploy Router-Tuning in half of the attention layers while maintaining the total MoD capacity at 50%. Across all models, Router-Tuning preserves performance compared to their original counterparts, demonstrating its effectiveness and adaptability across different architectures.
Figure 5: Ablation study on training samples.
Impact of number of MoD layers
In the main experiments, we deployed half of the layers with MoD. In Figure 4, we further explore the effect of the number of MoD layers. Our results indicate that applying MoD to up to half of the attention layers still maintain comparable performance. However, when further increasing the number of MoD layers, performance starts to degrade. We attribute this decline to the transformation of important shallow layers, which negatively impacts overall performance Men et al. (2024); He et al. (2024a, b). Therefore, preserving the density of shallow layers while applying MoD to deeper layers ensures its effectiveness.
Influence of Training Dataset
In Table 5, we next examine the impact of using different training datasets for Router-Tuning. We consider a variety of datasets, including Alpaca Taori et al. (2023), Evol-Instruct Xu et al. (2023), ShareGPT Zheng et al. (2023), and Llama-Pro Wu et al. (2024). Since Router-Tuning only fine-tunes the routers while keeping the backbone of the language models intact, changes in the training dataset do not significantly impact performance. However, Llama-Pro, which incorporates diverse training data from various domains, provides slightly better performance due to its broader data coverage.
@@
On the other hand, due to the small number of trainable parameters, Router-Tuning does not require a large amount of training samples. As shown in Figure 5, MoD layers are initially dense-activated and then sparsified. Although the initial sparsification steps lead to a drop in performance, subsequent Router-Tuning facilitates performance recovery. Notably, just 5K training samples are sufficient to effectively train the routers.
7 Router-Tuning Extensions
In this section, we explore the extensions of Router-Tuning to broader architectural designs and training paradigms. Specifically, we demonstrate its applicability to mainstream Mixture-of-Experts (MoE) architectures, and examine its integration with LoRA-based fine-tuning to jointly improve efficiency and downstream performance.
Table 6: Effectiveness of Router-Tuning integrated with LoRA fine-tuning, compared to using Router-Tuning and LoRA independently. The integrated approach achieves comparable performance while preserving efficiency, highlighting the compatibility and complementary strengths of the two methods.
@@ 7 Router-Tuning Extensions
The Expert’s redundancy in MoE has been widely demonstrated in recent works Lu et al. (2024); He et al. (2024a), e.g., the model still maintains comparable performance after removing certain layers. Therefore, we further extend Router-Tuning to MoE, where we take OLMoE Muennighoff et al. (2024) and DeepSeek-MoE Dai et al. (2024) as the backbones and equip each Expert network with Router-Tuning. Since these models deploy MoE at the token level, where each expert processes tokens from different sequences, we directly apply token-level Router-Tuning to these models. Here, we compare Router-Tuning with Expert Drop He et al. (2024a) that drops less important experts. For deployment, we configure an overall skipping ratio of 25%, achieved either by statically dropping 25% of layers or utilizing MoD on half of the MoE layers with 50% MoD capacity.
Figure 6: Expert-wise MoD capacity, visualizing the capacity utilization and selected frequency of individual experts within a specific layer.
In Table 4, while Expert Drop affects the performance significantly, Router-Tuning preserves more than 95% of the original performance, which demonstrates Router-Tuning’s effectiveness in the MoE architecture. In Figure 6, we further visualize the capacity utilization of individual experts within a specific layer, considering two perspectives for each expert: the frequency of expert selection (“Frequency”) and the MoD capacity allocated to each expert (“Capacity”).
Instead of removing a subset of experts (typically those being selected less frequently), Router-Tuning adopts a specialized and efficient utilization pattern, where high-frequency experts undergo greater sparsification compared to low-frequency experts. This approach retains the potential of all experts and fosters a more balanced utilization across the entire set of experts Fedus et al. (2022). Therefore, this approach maintains the potential of all experts and promotes a more balanced utilization across all experts.
@@ 8 Conclusion
Despite the progress we have made, our work still has limitations. First, while we have advanced MoD with Router-Tuning, other sophisticated training methods may exist that could further improve performance, warranting future exploration. Second, due to computational resource constraints, our experiments have been limited to a small set of models and tasks. Expanding this approach to other models and a broader range of tasks would be highly valuable for understanding its full potential.